# Comparison of the Migdal transition probabilities in electron-atom inelastic cross sections

### v1

@@ Abstract
The Migdal transition probabilities for dark matter scattering are compared to the total single-electron inelastic cross sections of electron-atom scattering for isolated Ar and Xe. The comparison is done by expressing the electron-atom scattering cross section by connecting the Migdal probability. The resultant errors are around $20\ \%$ for Ar but can be factors for Xe in $\sim$ 1 keV of incoming electron energy. The calculation is also compared with the cases of Coulomb potential and screened Coulomb potential.
@@ 1 Introduction
The energy eigenstates of atom are given in Sec. 2. The formalism uses the center of mass coordinates and is useful for treating long-range interactions. The interaction potential is derived in Sec. 3. The interactions are approximately described by the electron-electron interaction. Therefore, the shape of potential is the Coulomb one. The total excitation and ionization cross section is calculated in Sec. 4. Especially, the cross section is expressed by the Migdal transition factor of dark matter-atom scattering for single electron excitation/ionization with the dipole approximation given in Ref. Ibe:2017yqa . Then, the result is compared with a database value which is a combination of experimental and theoretical data. The calculation by using the screened Coulomb potential is shown in Sec. 5. The natural unit is used in this paper.
@@ 3 Interaction potential
Then, the invariant amplitude is related to the interaction potential $V(\vec{r}_{e,i}-\vec{r}_{\rm in})$ sakurai1967advanced by the Born approximation as Ibe:2017yqa
$V(\vec{r}_{e,i}-\vec{r}_{\rm in})=-\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}e^{i\vec{q}\cdot(\vec{r}_{e,i}-\vec{r}_{\rm in})}\frac{\mathcal{M}(q^{2})}{4m_{e}^{2}},$ (4)
where the invariant scattering amplitude is given as
$\mathcal{M}(q^{2})=$ $16\pi\alpha\frac{m_{e}^{2}}{q^{2}}.$ (5)
The variable $\vec{q}$ is the momentum transfer and $\vec{r}_{\rm in}$ is the coordinates of incoming electron. Remember that the $i$-th electron coordinate $\vec{r}_{e,i}$ is expressed by the relative coordinate $\vec{r}_{i}$ as $\vec{r}_{e,i}=\vec{r}_{i}+\vec{r}_{\rm CM}$.
@@ 4 Electron-atom inelastic scattering cross section
The initial state $|\Phi_{I}\rangle$ and final state $|\Phi_{F}\rangle$ are given by $|\Phi_{I/F}\rangle=\sqrt{2m_{A}}\sqrt{2m_{e}}|\Psi_{A}\rangle_{I/F}|\vec{p}_{\rm in}\rangle_{I/F}$. Then, the $T$-matrix by a potential $V(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})$ becomes
$iT_{FI}\approx$ $-2\pi i\delta(E_{F}-E_{I})\sum_{i=1}^{Z}\int d^{3}\vec{r}_{\rm in}d^{3}\vec{r}_{\rm CM}\left(\Pi_{j=1}^{Z}d^{3}\vec{r}_{j}\right)$
×$\times\Phi_{F}^{\dagger}(\vec{r}_{\rm in},\vec{r}_{\rm CM},\{\vec{r}_{k}\})V(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})\Phi_{I}(\vec{r}_{\rm in},\vec{r}_{\rm CM},\{\vec{r}_{l}\})$ (1)
$=$ $i(2\pi)^{4}\delta^{4}(p_{F}-p_{I})\frac{m_{A}}{m_{e}}\mathcal{M}(q^{2})Z_{FI}(\vec{q}),$ (2)
@@
The four-momentum in the laboratory frame is expressed by $p_{{\rm in},I}=(m_{e}+E_{{\rm kin},I},\vec{p}_{{\rm in},I})$, $p_{{\rm in},F}=(m_{e}+E_{{\rm kin},F},\vec{p}_{{\rm in},F})$, $p_{{\rm CM},I}=(m_{A},0)$, and $p_{{\rm CM},F}=(E_{A,F},\vec{p}_{A,F})$, where $E_{\rm kin}\approx\vec{p}_{\rm in}\,^{2}/(2m_{e})$ is the kinetic energy of incoming electron in the laboratory frame. The momentum of final state atom becomes $\vec{p}_{A,F}\approx\vec{p}_{n,F}+\vec{p}_{i,F}$. Here, $\vec{p}_{i,F}=0$ for excitation. Then, $m_{A}=m_{n}+Zm_{e}+E_{{\rm re},I}$ and $E_{A,F}\approx m_{n}+Zm_{e}+E_{{\rm re},F}+\vec{p}_{n,F}\,^{2}/(2m_{n})$ are followed. Finally, the minimum and maximum momentum transfer, $q^{2}_{\rm min}\approx 2m_{e}\Delta E$ and $q^{2}_{\rm max}=2m_{e}E_{{\rm kin},I}$, are obtained, where $\Delta E=E_{e,F}-E_{e,I}>0$ is the transition energy of electron in the atom.
@@
$\sigma(\vec{p}_{{\rm in},I}\,^{2})\approx$ $\sum_{F}\int\frac{d^{3}\vec{p}_{{\rm in},F}}{(2\pi)^{3}2p_{{\rm in},F}^{0}}\frac{d^{3}\vec{p}_{{\rm CM},F}}{(2\pi)^{3}2p_{{\rm CM},F}^{0}}\frac{1}{4\sqrt{(p_{{\rm CM},I}\cdot p_{{\rm in},I})^{2}-m_{A}^{2}m_{e}^{2}}}$
×$\times\left(\frac{m_{A}}{m_{e}}\right)^{2}|\mathcal{M}(q^{2})|^{2}|Z_{FI}(\vec{q})|^{2}(2\pi)^{4}\delta^{4}(p_{{\rm in},F}+p_{{\rm CM},F}-p_{{\rm in},I}-p_{{\rm CM},I})$ (7)
$\approx$ $\frac{4\pi\alpha^{2}m_{e}^{2}}{\vec{p}_{{\rm in},I}\,^{2}}\sum_{F}\left|z_{FI}\right|^{2}\log\frac{q^{2}_{\rm max}}{q^{2}_{\rm min}},$ (8)
@@
$\sigma(E_{{\rm kin},I})=\frac{2\pi\alpha^{2}m_{e}}{E_{{\rm kin},I}}\sum_{F}\left|z_{FI}\right|^{2}\log\frac{E_{{\rm kin},I}}{\Delta E}.$ (9)
In the cross section, the possible final states $F$ are summed. In Eq. (8), the average/sum of magnetic quantum number for the initial/final states are taken, which makes the factor $\left|z_{FI}(\theta_{qr})\right|^{2}$ isotropic for the angular direction $\left|z_{FI}(\theta_{qr})\right|^{2}\to\left|z_{FI}\right|^{2}$ Ibe:2017yqa . Note that the sum can include the integration of ionized electron energy.
@@
Figure 1: The total excitation and ionization cross section of electron-atom scattering $\sigma(E_{{\rm kin},I})$ as the function of kinetic energy for incoming electron in laboratory frame $E_{{\rm kin},I}$. The calculation results using the transition probabilities of Ref. Ibe:2017yqa are shown by the solid lines and the comparison values are shown by the dashed lines. The red lines show the single excitation case and blue lines show the single ionization case. This calculation uses the Coulomb potential. (top:) For Ar case. The database value for comparison is taken from Phelps Phelps:2024 . (bottom:) For Xe case. The database value for comparison is taken from Morgan Morgan:2024 . The ionization data is limited up to 1 keV.
@@
The calculation result of $\sigma(E_{{\rm kin},I})$ is shown in Fig. 1. For Ar, the calculation error of ionization is around $20\ \%$ from the database value in 300 eV $\lesssim E_{{\rm kin},I}<10$ keV. That of excitation has local maximum of around $60\ \%$ at 100 eV and gradually decreasing by increasing $E_{{\rm kin},I}$. For Xe, the error of ionization is around 60 % at 1 keV and that of excitation is roughly factor 3 at 1 keV. This error dependency is consistent with that the Born approximation is a good approximation in the keV energy region bonham1974high ; garcia1999electron . Also, the overestimation near the threshold energy is consistent with Ref. ariyasinghe2004total which compares to the Bethe-Born theory inokuti1971inelastic . Note that the probability is dominated by the excitation/ionization from the outermost shell, i.e., $3p$ states for Ar and $5p$ states for Xe.
5 Screened Coulomb potential
The overestimation near threshold energy will come from the use of the Coulomb potential for the interaction potential. The actual potential will be the screened Coulomb potential which is obtained by introducing the screening mass term from the electron number density in Eq. (1). Here, the mass terms are assumed to come from the inner shell electrons because the dominant contribution of excitation and ionization comes from the outermost shell. Therefore, the Maxwell equation becomes
$(\nabla^{2}-r_{0}^{-2})\phi(\vec{r})=-\left[Ze\delta(\vec{r}-\vec{r}_{n})-\sum_{\rm outermost}e\delta(\vec{r}-\vec{r}_{e,i})\right],$ (1)
where $r_{0}$ is the screening radius, additional parameter. Then, the screened potential is obtained as
$\phi(\vec{r})=$ $\frac{Ze}{4\pi|\vec{r}-\vec{r}_{n}|}e^{-\frac{|\vec{r}-\vec{r}_{n}|}{r_{0}}}-\sum_{\rm outermost}\frac{e}{4\pi|\vec{r}-\vec{r}_{e,i}|}e^{-\frac{|\vec{r}-\vec{r}_{e,i}|}{r_{0}}}$ (2)
$=$ $\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}\left(\frac{Ze}{q^{2}+r_{0}^{-2}}e^{-i\vec{q}\cdot(\vec{r}-\vec{r}_{n})}-\sum_{\rm outermost}\frac{e}{q^{2}+r_{0}^{-2}}e^{-i\vec{q}\cdot(\vec{r}-\vec{r}_{e,i})}\right).$ (3)
The cross section becomes
$\sigma(E_{{\rm kin},I})=$ $\frac{2\pi\alpha^{2}m_{e}}{E_{{\rm kin},I}}\sum_{F}\left|z_{FI}\right|^{2}$
×$\times\left(\frac{1}{1+2r_{0}^{2}m_{e}E_{{\rm kin},I}}-\frac{1}{1+2r_{0}^{2}m_{e}\Delta E}+\log\frac{1+2r_{0}^{2}m_{e}E_{{\rm kin},I}}{1+2r_{0}^{2}m_{e}\Delta E}\right).$ (4)
Figure 2: The excitation and ionization cross section of outermost shell electrons $\sigma(E_{{\rm kin},I})$ as the function of kinetic energy for incoming electron in laboratory frame $E_{{\rm kin},I}$. The calculation results using the transition probabilities of Ref. Ibe:2017yqa are shown by the solid lines and the comparison values are shown by the dashed lines. The red lines show the single excitation case and blue lines show the single ionization case. This calculation uses the screened Coulomb potential. (top:) For Ar case. The atomic radius $r_{0}\approx 0.7$ clementi1967atomic is used. The only $3p$ state electrons are contributed. The database value for comparison is taken from Phelps Phelps:2024 . (bottom:) For Xe case. The atomic radius $r_{0}\approx 1.1$ clementi1967atomic is used. The only $5p$ state electrons are contributed. The database value for comparison is taken from Morgan Morgan:2024 . The ionization data is limited up to 1 keV.
The result for the screened Coulomb potential is shown in Fig. 2. The errors become smaller near threshold region. As a result, it becomes better for excitation, but, it deviates for ionization at higher energy region. The deviation becomes around $50\ \%$ for Ar ionization at 10 keV.

### v3

@@ Abstract
The Migdal transition probabilities for dark matter scattering are compared to the total single-electron inelastic cross sections of electron-atom scattering for isolated Ar and Xe. The comparison is done by expressing the electron-atom scattering cross section by connecting the Migdal probability in the Born approximation. The resultant errors are around $20\ \%$ for Ar and $60\ \%$ for Xe in $\sim$ 1 keV of incoming electron energy. The transition is dominated by the outermost shell electrons. The error for Xe atom will become better for high energy because of the approximation and lack of data.
@@ 1 Introduction
The energy eigenstates of atom are given in Sec. 2. The formalism uses the center of mass coordinates and is useful for treating long-range interactions. The interaction potential is derived in Sec. 3. The interactions are approximately described by the electron-electron interaction. Therefore, the shape of potential is the Coulomb one. The total excitation and ionization cross section is calculated in Sec. 4. Especially, the cross section is expressed by the Migdal transition factor of dark matter-atom scattering for single electron excitation/ionization with the dipole approximation given in Ref. Ibe:2017yqa . Then, the result is compared with a database value which is a combination of experimental and theoretical data. The natural unit is used in this paper.
@@ 3 Interaction potential
Then, the interaction Lagrangian for the ionization scattering of electron and atom is given by
$\mathcal{L}_{\rm int}\approx e\bar{\psi}_{e}\gamma^{\mu}\psi_{e}A_{\mu},$ (4)
where $\psi_{e}$ and $A_{\mu}$ are the electron and photon fields and $\gamma^{\mu}$ is the gamma matrix. The invariant scattering amplitude is given by the $t$-channel and $u$-channel processes as
$\mathcal{M}_{i}\approx 16\pi\alpha m_{e}^{2}\left(\frac{\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\delta_{s_{i}^{F}s_{i}^{I}}}{t_{i}}-\frac{\delta_{s_{i}^{F}s_{\rm in}^{I}}\delta_{s_{\rm in}^{F}s_{i}^{I}}}{u_{i}}\right)$ (5)
under the non-relativistic approximation, where $t_{i}=(p_{{\rm in},I}-p_{{\rm in},F})^{2}$ and $u_{i}=(p_{{\rm in},I}-p_{e,i,F})^{2}$ are the Mandelstam variables for $i$-th electron in the atom. The subscript $I$ and $F$ represent the label for initial and final states, $s$ represents the spin, and an incoming electron is labeled by in.
Then, the invariant amplitude is related to the interaction potential $V(\vec{r}_{e,i}-\vec{r}_{\rm in})$ sakurai1967advanced by the Born approximation. The potential is written as
$V(\vec{r}_{e,i}-\vec{r}_{\rm in})=V_{t}(\vec{r}_{e,i}-\vec{r}_{\rm in})+V_{u}(\vec{r}_{e,i}-\vec{r}_{\rm in}),$ (6)
where Ibe:2017yqa
$V_{t}(\vec{r}_{e,i}-\vec{r}_{\rm in})=$ $-\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}e^{i\vec{q}\cdot(\vec{r}_{e,i}-\vec{r}_{\rm in})}\frac{\mathcal{M}_{t}(q^{2})}{4m_{e}^{2}},$ (7)
$V_{u}(\vec{r}_{e,i}-\vec{r}_{\rm in})=$ $-\int\frac{d^{3}\vec{q}_{u}}{(2\pi)^{3}}e^{i\vec{q}_{u}\cdot(\vec{r}_{e,i}-\vec{r}_{\rm in})}\frac{\mathcal{M}_{u}(q_{u}^{2})}{4m_{e}^{2}}$ (8)
and the invariant scattering amplitude is given as
$\mathcal{M}_{t}(q^{2})=$ $-16\pi\alpha\frac{m_{e}^{2}}{q^{2}},$ (9)
$\mathcal{M}_{u}(q_{u}^{2})=$ $16\pi\alpha\frac{m_{e}^{2}}{q_{u}^{2}}.$ (10)
Here, $s_{i}=(p_{{\rm in},I}+p_{e,i,I})^{2}$ is the Mandelstam variable, $\vec{q}_{u}=\vec{p}_{{\rm in},I}-\vec{p}_{e,i,F}$ is the $u$-channel momentum transfer, $\vec{q}=\vec{p}_{{\rm in},I}-\vec{p}_{{\rm in},F}$ is the momentum transfer of $t$-channel, and $\vec{r}_{\rm in}$ is the coordinates of incoming electron. Remember that the $i$-th electron coordinate $\vec{r}_{e,i}$ is expressed by the relative coordinate $\vec{r}_{i}$ as $\vec{r}_{e,i}=\vec{r}_{i}+\vec{r}_{\rm CM}$.
@@ 4 Electron-atom inelastic scattering cross section
The initial state $|\Phi_{I}\rangle$ and final state $|\Phi_{F}\rangle$ for asymptotic states are given by $|\Phi_{I/F}\rangle=\sqrt{2m_{A}}\sqrt{2m_{e}}|\Psi_{A}\rangle_{I/F}|\vec{p}_{\rm in}\rangle_{I/F}$. The incoming electron states can be taken as plane waves because the atom is regarded to be neutral far from the atom even the atom is ionized. Then, the $T$-matrix by a $t$-channel potential $V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})$ becomes
$iT_{t,FI}=$ $-2\pi i\delta(E_{F}-E_{I})\sum_{i=1}^{Z}\int d^{3}\vec{r}_{\rm in}d^{3}\vec{r}_{\rm CM}\left(\Pi_{j=1}^{Z}d^{3}\vec{r}_{j}\right)$
×$\times\Phi_{F}^{\dagger}(\vec{r}_{\rm in},\vec{r}_{\rm CM},\{\vec{r}_{k}\})V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})\Phi_{I}(\vec{r}_{\rm in},\vec{r}_{\rm CM},\{\vec{r}_{l}\})$ (1)
$=$ $i(2\pi)^{4}\delta^{4}(p_{F}-p_{I})\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\frac{m_{A}}{m_{e}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q}),$ (2)
@@
The $T$-matrix for $u$-channel potential becomes
$iT_{u,FI}=$ $i(2\pi)^{4}\delta(E_{F}-E_{I})\frac{m_{A}}{m_{e}}\int\frac{d^{3}\vec{q}_{u}}{(2\pi)^{3}}\delta^{3}(\vec{q}_{u}+\vec{p}_{{\rm CM},I}-\vec{p}_{{\rm CM},u,F})\mathcal{M}_{u}(q_{u}^{2})$
×$\times\int d^{3}\vec{r}_{\rm in}\phi_{E_{e},F}^{\dagger}(\vec{r}_{\rm in})e^{-i\vec{q}_{u}\cdot\vec{r}_{\rm in}}\langle\vec{r}_{\rm in}|\vec{p}_{{\rm in},I}\rangle$
×$\times\int d^{3}\vec{r}\langle\vec{p}_{{\rm in},F}|\vec{r}\rangle\phi_{E_{e},I}(\vec{r})e^{i\vec{q}_{u}\cdot\vec{r}}.$ (7)
Here, $\vec{p}_{{\rm CM},u,F}=\vec{p}_{{\rm CM},F}-\vec{p}_{i,F}+\vec{p}_{{\rm in},F}$ because of the $u$-channel process. The integrations by $\vec{r}_{\rm in}$ and $\vec{r}$ are actually the Fourier transformations of the projected single-electron wave functions. Assuming that the plane wave $\langle\vec{r}|\vec{p}\rangle$ is approximately equal to the Coulomb wave $\phi(\vec{r})$ (see Ref. gozem2015photoelectron for the discussion of the differences), the above equation becomes
$iT_{u,FI}\approx$ $i(2\pi)^{4}\delta(E_{F}-E_{I})\delta^{3}(\vec{p}_{{\rm in},I}+\vec{p}_{{\rm CM},I}-\vec{p}_{{\rm in},F}-\vec{p}_{{\rm CM},F})\delta_{s_{i}^{F}s_{\rm in}^{I}}\frac{m_{A}}{m_{e}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u}).$ (8)
Here, the transition matrix is given as
$Z_{{\rm in}I}(\vec{q}_{u})=\int d^{3}\vec{r}\phi_{{\rm in},F}^{\dagger}(\vec{r})e^{i\vec{q}_{u}\cdot\vec{r}}\phi_{E_{e},I}(\vec{r}),$ (9)
where $\phi_{{\rm in},F}(\vec{r})$ is the Coulomb wave function with energy for incoming electron and $\vec{q}_{u}=\vec{p}_{{\rm in},I}-\vec{p}_{i,F}=\vec{p}_{{\rm CM},F}-\vec{p}_{{\rm CM},I}-\vec{p}_{i,F}+\vec{p}_{{\rm in},F}$ is again the momentum transfer for $u$-channel process. This approximation will be not good for excitation processes because the excitation states are actually bound states.
Finally, the $T$-matrix is obtained as
$iT_{FI}=$ $iT_{t,FI}+iT_{u,FI}$ (10)
$=$ $i(2\pi)^{4}\delta^{4}(p_{F}-p_{I})\frac{m_{A}}{m_{e}}\left(\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q})+\delta_{s_{i}^{F}s_{\rm in}^{I}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u})\right).$ (11)
The four-momentum in the laboratory frame is expressed by $p_{{\rm in},I}=(m_{e}+E_{{\rm kin},I},\vec{p}_{{\rm in},I})$, $p_{{\rm in},F}=(m_{e}+E_{{\rm kin},F},\vec{p}_{{\rm in},F})$, $p_{{\rm CM},I}=(m_{A},0)$, and $p_{{\rm CM},F}=(E_{A,F},\vec{p}_{A,F})$, where $E_{\rm kin}\approx\vec{p}_{\rm in}\,^{2}/(2m_{e})$ is the kinetic energy of incoming electron in the laboratory frame. The momentum of final state atom becomes $\vec{p}_{A,F}\approx\vec{p}_{n,F}+\vec{p}_{i,F}$. Here, $\vec{p}_{i,F}=0$ for excitation. Then, $m_{A}=m_{n}+Zm_{e}+E_{{\rm re},I}$ and $E_{A,F}\approx m_{n}+Zm_{e}+E_{{\rm re},F}+\vec{p}_{n,F}\,^{2}/(2m_{n})$ are followed. Finally, the minimum and maximum momentum transfer, $q^{2}_{\rm min}\approx 2m_{e}\Delta E$ and $q^{2}_{\rm max}\approx 2m_{e}(E_{{\rm kin},I}-\Delta E)$, are obtained, where $\Delta E=E_{e,F}-E_{e,I}>0$ is the transition energy of electron in the atom. The $u$-channel momentum transfer becomes $\vec{q}_{u}\,^{2}\approx 2m_{e}E_{{\rm kin},I}-\vec{q}\,^{2}$.
@@
$\sigma(\vec{p}_{{\rm in},I}\,^{2})=$ $\frac{1}{2!}\frac{1}{2}\sum_{F,s_{\rm in}^{F/I}}\int\frac{d^{3}\vec{p}_{{\rm in},F}}{(2\pi)^{3}2p_{{\rm in},F}^{0}}\frac{d^{3}\vec{p}_{{\rm CM},F}}{(2\pi)^{3}2p_{{\rm CM},F}^{0}}\frac{1}{4\sqrt{(p_{{\rm CM},I}\cdot p_{{\rm in},I})^{2}-m_{A}^{2}m_{e}^{2}}}$
×$\times\left(\frac{m_{A}}{m_{e}}\right)^{2}(2\pi)^{4}\delta^{4}(p_{{\rm in},F}+p_{{\rm CM},F}-p_{{\rm in},I}-p_{{\rm CM},I})$
×$\times\left|\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q})+\delta_{s_{i}^{F}s_{\rm in}^{I}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u})\right|^{2}.$ (12)
In the cross section, the possible final states $F$ and spins are summed or averaged. The electron spins of final states $s_{i}^{F}$ are included in the label $F$. Note that the sum includes the integration of ionized electron energy. This equation is reduced as
$\sigma(\vec{p}_{{\rm in},I}\,^{2})=$ $\frac{2\pi\alpha^{2}m_{e}^{2}}{\vec{p}_{{\rm in},I}\,^{2}}\sum_{F}\left[\left|z_{FI}\right|^{2}\log\frac{q^{2}_{\rm max}}{q^{2}_{\rm min}}+\left|z_{{\rm in}I}\right|^{2}\log\frac{2m_{e}E_{{\rm kin},I}-q_{\rm min}^{2}}{2m_{e}E_{{\rm kin},I}-q_{\rm max}^{2}}\right.$
$\left.-2{\rm Re}\left[z_{FI}z_{{\rm in}I}^{*}\right]\left(\sin^{-1}\sqrt{\frac{q_{\rm max}^{2}}{2m_{e}E_{{\rm kin},I}}}-\sin^{-1}\sqrt{\frac{q_{\rm min}^{2}}{2m_{e}E_{{\rm kin},I}}}\right)\right],$ (13)
@@
$\sigma(E_{{\rm kin},I})=$ $\frac{2\pi\alpha^{2}m_{e}}{E_{{\rm kin},I}}\sum_{F}\left[\left|z_{FI}\right|^{2}\log\frac{E_{{\rm kin},I}-\Delta E}{\Delta E}\right.$
$\left.-{\rm Re}\left[z_{FI}z_{{\rm in}I}^{*}\right]\left(\sin^{-1}\sqrt{1-\frac{\Delta E}{E_{{\rm kin},I}}}-\sin^{-1}\sqrt{\frac{\Delta E}{E_{{\rm kin},I}}}\right)\right].$ (14)
To obtain Eq. (13), the average/sum of magnetic quantum number for the initial/final states are taken, which makes the factor $\left|z_{FI}(\theta_{qr})\right|^{2}$ isotropic for the angular direction $\left|z_{FI}(\theta_{qr})\right|^{2}\to\left|z_{FI}\right|^{2}$ Ibe:2017yqa . Also, the fact that the transition matrix element does not change the spins is used. In Eq. (14), the relation $\sum_{F}\left|z_{FI}\right|^{2}=\sum_{F}\left|z_{{\rm in}I}\right|^{2}$ is used. The final states of $z_{FI}z_{{\rm in}I}^{*}$ are related by the energy conservation. The first term is actually the cross section by the Coulomb potential without the identification factor of the interacting electrons $1/2!$.
@@
Figure 1: The total excitation and ionization cross section of electron-atom scattering $\sigma(E_{{\rm kin},I})$ as the function of kinetic energy for incoming electron in laboratory frame $E_{{\rm kin},I}$. The calculation results using the transition probabilities of Ref. Ibe:2017yqa are shown by the solid lines and the comparison values are shown by the dashed lines. The red lines show the single excitation case and blue lines show the single ionization case. (top:) For Ar case. The database value for comparison is taken from Phelps Phelps:2024 . (bottom:) For Xe case. The database value for comparison is taken from Morgan Morgan:2024 . The ionization data is limited up to 1 keV.
@@
The calculation result of $\sigma(E_{{\rm kin},I})$ is shown in Fig. 1. For Ar, the calculation error of ionization is around $20\ \%$ from the database value in 300 eV $\lesssim E_{{\rm kin},I}<10$ keV. That of excitation has local maximum of around $60\ \%$ at 100 eV and gradually decreasing by increasing $E_{{\rm kin},I}$. For Xe, the error of ionization is around 60 % at 1 keV and that of excitation is roughly factor 3 at 1 keV. This error dependency is consistent with that the Born approximation is a good approximation in the keV energy region bonham1974high ; garcia1999electron . The interference term decreases the cross section for low energy and negligible for high energy. Neglected the interference term, the overestimation near the threshold energy is consistent with Ref. ariyasinghe2004total which compares to the Bethe-Born theory inokuti1971inelastic . Note that the probability is dominated by the excitation/ionization from the outermost shell, i.e., $3p$ states for Ar and $5p$ states for Xe.